I need help figuring out the next number in this sequence. the problem is... 3;9;81;6;561
Answer by checkley77(12569)

IFYOUR PROBLEM IS 3,9,81 &,6,561 THEN:
3^2=9
9^2=81
81^2=6,561
6,561^2=43,046,721